A Christmas light set consists of 40 identical bul
Answer #1

(8) (a) is correct.

In the series circuit, any current that flows through one bulb must go through the other bulbs as well, so each bulb draws the same current. Since all 40 bulbs have the same resistance, the voltage drop across each one is the same and equals (1/40)th of the applied voltage, or 120/40 = 3 volts.

Now, total power = 140 W implies,

40\times I^{2} R=140

\Rightarrow 40\times \left (\frac{V}{40R} \right )^{2}\times R=140

\Rightarrow \frac{120^{2}}{40R}=140

\Rightarrow \boldsymbol{R=2.57\Omega}

(9) Net charge is given by:

q_{net}=n_{p}e-n_{e}e=(8.67\times10^{6}-6.51\times10^{6})\times1.602\times10^{-19}\ C=\boldsymbol{3.46\times10^{-13}\ C}

Thus, (b) is correct.

(10) Force on proton = Weight of proton (W) = mpg = (1.6726219 × 10-27 x 9.81) N = 1.64 x 10-26 N

Let x be the required distance. Then,

F_{e}=W

\Rightarrow \frac{k_{e}qe}{x^{2}}=W

\Rightarrow \frac{9\times10^{9}\times0.82\times10^{-9}\times1.602\times10^{-19}}{x^{2}}=1.64\times10^{-26}

Solving, we get x = 8480 m.

Thus (a) is correct.

(13) Electric field due to a point charge is given by:

E_{5.01\mu C}=\frac{k_{e}q}{r^{2}}=\frac{9\times10^{9}\times5.01\times10^{-6}}{1.65^{2}}\approx \boldsymbol{16500\ N/C}

Thus (a) is correct.

(14) The voltage difference between the two plates can be expressed in terms of the work done on a positive test charge q when it moves from the positive to the negative plate.

V=\frac{work\ done}{charge}=\frac{Fd}{q}=Ed=3.2\times10^{4}\times0.58\times10^{-3}\ V=\boldsymbol{18.6\ V}

Thus (a) is correct.

(15) Charge on capacitor plate is related to potential difference by following relation -

Q=CV

Therefore,

Q_{plate}=0.49\times10^{-6}\times12=5.88\times10^{-6}C=\boldsymbol{5.88\mu C}

Thus (b) is correct.